A basketball player throws the ball at the different angles. At which of the following angles will the ball achieve the maximum height? A. 35.0° B. 45.0° C. 65.0° D. 70.0°


1
Expert's answer
2022-04-16T15:14:30-0400

Maximum height can be calculated by the formula:

"h_{max}=\frac{v_0^2\sin^2\alpha}{2g}" , where "v_0" - initial velocity, "g=9.81" "m\/s^2" .

"\sin \alpha" grows if "\alpha \in (0; \frac{\pi}{2})" so we can conclude that the ball will achieve the maximum height if "\alpha =70.0 \degree" .